OpenStudy (anonymous):

Solve the logarithmic equation. Be sure to reject any value that is not in the domain of the original logarithmic expressions. Give the exact answer. ln (x - 6) + ln (x + 1) = ln (x - 15) A. {3, -3} B. {-3} D. Ø

hartnn (hartnn):

what have you tried?

hartnn (hartnn):

log A+log B = log AB

OpenStudy (anonymous):

ln A+ln B=ln C implies ln AB=ln C that means AB=C . you have to solve this equation and take those points for which all the expressions in the logarithms are greater than 0. @SkiTTleoooo47

OpenStudy (anonymous):

(x - 6)(x + 1)=(x - 15) gives x^2-5x-6=x-15 this implies x^2-6x+9=0 which implies (x-3)^2=0 which implies x=3,3 but see ln(x-6)=ln(3-6)=ln(-3) is invalid . Hence the answer is option D.@SkiTTleoooo47

OpenStudy (anonymous):

Oh , thanks you :)